# The Nesterov–Spokoiny Acceleration Achieves Strict o(1/k2) Convergence

### v4

@@ 3 The NSA Algorithm with Inexact Gradient Oracle
Recently, there has been a growing demand for zeroth-order optimization techniques, primarily driven by the need for fine-tuning in contemporary deep learning (e.g., Zhang et al., (2025)). Additionally, many learning scenarios provide access solely to zeroth-order information of the objective function Larson et al., (2019). In such cases, we can estimate the gradient first Flaxman et al., (2005); Nesterov and Spokoiny, (2017); Feng and Wang, (2023); Wang, (2022) and then apply gradient algorithms to solve the problem. Such gradient estimators can be abstractly summarized as an inexact gradient oracle, defined as follows.
@@
Such gradient oracle $\mathcal{G}$ can be guaranteed by zeroth-order information. Formal statements for these properties are in Theorems 2 and 3. Theorem 2 has been proved in the literature by different authors Flaxman et al., (2005); Nesterov and Spokoiny, (2017); Feng and Wang, (2023). The key message is that inexact gradient oracles that satisfy conditions (H1) and (H2) can be derived using zeroth-order information. A simple method that ensures conditions (H1) and (H2) is performing finite differences along each coordinate to derive the gradient estimator. Specifically, we consider the most straightforward finite difference gradient estimator defined as follows:
@@
where $\mathbf{e}_{i}$ is the $i$-th unit vector in a fixed canonical basis. The straightforward estimator in (9) serves as an example that fulfills conditions (H1) and (H2). It is important to note that this estimator is entirely deterministic, making $\xi$ a mere placeholder, and thus $\mathbb{E}_{\xi}$ acts as an identity operator. However, (9) is not the only simple estimator that meets these criteria; the estimator presented in Feng and Wang, (2023) also satisfies conditions (H1) and (H2).
Theorem 2 (Flaxman et al., (2005); Nesterov and Spokoiny, (2017); Feng and Wang, (2023)).
@@
Consider a function $f\in\mathscr{F}$, where $\mathscr{F}$ denotes Nesterov’s $\mathscr{F}$ functions Nesterov, (2003, 2018). If at any $\mathbf{x}\in\mathbb{R}^{n}$, we can access the function value $f(\mathbf{x})$, then there exists an oracle $\mathcal{G}$ such that (1) $\mathcal{G}$ satisfies property (H2), and (2) $\mathcal{G}$ can be explicitly calculated using zeroth-order information.
@@
7:   $\mathbf{x}_{k+1}^{\prime}=\mathbf{y}_{k}-2\eta\mathbf{g}_{k}(\mathbf{y}_{k})$. /* $\eta$ is the step size. */
8:   $\mathbf{x}_{k+1}^{\prime\prime}=\mathbf{x}_{k}-2\eta\mathbf{g}_{k}(\mathbf{x}_{k})$.
9:   $\mathbf{x}_{k+1}=\begin{cases}\mathbf{x}_{k+1}^{\prime},&\text{ if }f(\mathbf{x}_{k+1}^{\prime})\leq f(\mathbf{x}_{k+1}^{\prime\prime})\\ \mathbf{x}_{k+1}^{\prime\prime},&\text{ otherwise. }\end{cases}$
10:   $\mathbf{z}_{k+1}^{\prime}=\mathbf{z}_{k}-\frac{\eta}{\alpha_{k}}\mathbf{g}_{k}(\mathbf{y}_{k})$.
11:   $\begin{cases}\text{{Option I: }}\mathbf{z}_{k+1}=\mathbf{z}_{k+1}^{\prime}.\\ \text{{Option II: }}\mathbf{z}_{k+1}=\Pi_{\mathbb{B}_{\mu}}(\mathbf{z}_{k+1}^{\prime}),\text{where $\Pi_{\mathbb{B}_{\mu}}$ is the projection to the ball ofradius $\mu$},\end{cases}$
@@
$\liminf_{k\to\infty}\beta^{k}\|f(\mathbf{x}_{k})\|=0\;\text{ for some $\beta>1$},$ (12)
@@
Let $f\in\mathscr{F}_{L}^{1,1}(\mathbb{R}^{n})$ and let $\eta\in(0,\frac{1}{2L}]$. Let $\alpha_{k}=\frac{p}{k+p}$ for some $p>0$. Let $\mathbf{x}_{k}$ be governed by the NSA algorithm. If the gradient estimation oracle $\mathcal{G}$ satisfies properties (H1) and (H2), then there exists an error sequence $\{\epsilon_{k}\}_{k\in\mathbb{N}}$, such that $\sum_{k=1}^{\infty}\frac{\epsilon_{k}}{\alpha_{k}}<\infty$, and for all $k=0,1,2,\cdots$, it holds that
@@
$\frac{3}{4}\|\mathcal{G}(f,\mathbf{y}_{k},\epsilon^{\prime};\xi_{k}^{\prime})\|^{2}\leq\nabla f(\mathbf{x})^{\top}\mathcal{G}(f,\mathbf{y}_{k},\epsilon^{\prime};\xi_{k}^{\prime})\leq\frac{5}{4}\|\mathcal{G}(f,\mathbf{y}_{k},\epsilon^{\prime};\xi_{k}^{\prime})\|^{2},$ (13)
@@
$\frac{3}{4}\|\mathbf{g}_{k}(\mathbf{y}_{k})\|^{2}\leq\nabla f(\mathbf{x})^{\top}\mathbf{g}_{k}(\mathbf{y}_{k})\leq\frac{5}{4}\|\mathbf{g}_{k}(\mathbf{y}_{k})\|^{2},\quad a.s.$
@@
$\leq$ $\;f(\mathbf{y}_{k})+\left(-2\eta+2L\eta^{2}\right)\|\mathbf{g}_{k}(\mathbf{y}_{k})\|^{2},$
where the last inequality follows from (13). Since $\eta\in(0,\frac{1}{2L}]$, we know
@@
$f(\mathbf{x}_{k+1}^{\prime\prime})\leq f(\mathbf{y}_{k})+\nabla f(\mathbf{y}_{k})^{\top}\left(\mathbf{x}_{k+1}^{\prime\prime}-\mathbf{y}_{k}\right)+\frac{L}{2}\|\mathbf{x}_{k+1}^{\prime\prime}-\mathbf{y}_{k}\|^{2}\leq f(\mathbf{x}_{k})-\eta\|\mathbf{g}_{k}(\mathbf{y}_{k})\|^{2}.$ (15)
@@
$\leq$ $\;\sum_{\kappa\in\mathbb{Q}_{+}}\mathbb{P}\left(\liminf_{k\to\infty}k^{2}\log k\log\log k(f(\mathbf{x}_{k})-f^{*})\geq\kappa\right)$
[…]
@@ 4 NSA for Nonsmooth Objective
[…]
@@
$\begin{split}\mathbf{x}_{k+1}=&\;\text{{prox}}_{s,f}\left((1-\alpha_{k})\mathbf{x}_{k}+\alpha_{k}\mathbf{z}_{k}\right)\\ \mathbf{z}_{k+1}=&\;\mathbf{z}_{k}+\frac{\eta}{\alpha_{k}}\left(\mathbf{x}_{k+1}-\big{(}(1-\alpha_{k})\mathbf{x}_{k}+\alpha_{k}\mathbf{z}_{k}\big{)}\right),\end{split}$ (21)
@@
Let $\partial f(\mathbf{x})$ be the set of subgradients of $f$ at $\mathbf{x}$. By the optimality condition for subgradients (e.g., Rockafellar and Wets, (1998)), there exists $\mathbf{g}_{k+1}\in\partial f(\mathbf{x}_{k+1})$, such that
@@
$=$ $\;\frac{\gamma_{k+1}}{2}\left\|\mathbf{z}_{k}-\mathbf{x}^{*}\right\|^{2}+\left<\mathbf{g}_{k+1},\alpha_{k}\mathbf{x}^{*}+(1-\mathbf{x}_{k})\mathbf{x}_{k}-s\mathbf{g}_{k+1}\right>$
@@
$\leq$ $\;\frac{\gamma_{k+1}}{2}\left\|\mathbf{z}_{k}-\mathbf{x}^{*}\right\|^{2}+\left<\mathbf{g}_{k+1},\alpha_{k}\mathbf{x}^{*}+(1-\mathbf{x}_{k})\mathbf{x}_{k}\right>+f(\mathbf{x}_{k+1})-f(\mathbf{x}^{*})$ (since $s\geq\frac{\eta}{2}$)
@@
where the last two lines use convexity of the function $f$. Rearranging terms gives that $\frac{\Delta_{k+1}}{\gamma_{k+1}}\leq\frac{\Delta_{k+1}}{\gamma_{k}}$. Therefore,
$\frac{\Delta_{k+1}}{\gamma_{k+1}}\leq\frac{\Delta_{k+1}}{\gamma_{k}}\cdots\leq\frac{\Delta_{1}}{\gamma_{1}},$
@@ 4.1 A Continuous-time Perspective
In this section, we provide a continuous-time analysis for a proximal version of the NSA algorithm (21). In order for us to find a proper continuous-time trajectory of the discrete-time algorithm, Taylor approximation in Hilbert spaces will be used. Therefore, we want the objective function to be smooth. Otherwise, the error terms can get out-of-control and the correspondence between the discrete-time algorithm and continuous-time trajectory may break. After an ODE is derived for the smooth case via Taylor approximation, we could analogously write out an differential inclusion for the nonsmooth case, with derivatives defined in a subdifferential sense (e.g., Attouch and Peypouquet, (2016)).
@@
Let $\{\mathbf{x}_{k}\}_{k\in\mathbb{N}}$ and $\{\mathbf{z}_{k}\}_{k\in\mathbb{N}}$ be associated with two curves $X(t)$ and $Z(t)$ ($t\geq 0$). For the first equation in (22), we introduce a correspondence between discrete-time index $k$ and continuous time index $t$: $t=k\sqrt{\eta}$. With this we have $X(t)\approx\mathbf{x}_{k}$, $X(t+\sqrt{\eta})\approx\mathbf{x}_{k+1}$, and $\alpha_{k}=\frac{p}{k+p}=\frac{\sqrt{\eta}p}{t+\sqrt{\eta}p}=\frac{\sqrt{\eta}p}{t}+o(\sqrt{\eta})$. Thus, it holds that
@@
and when $s=o(1)$, we Taylor approximate the trajectories in the Hilbert space $\mathcal{H}$ Cartan et al., (2017)
@@
[…]
@@ 4.1.1 A Single Second-order ODE for (22)
Note that the single second-order ODE recovers the high-resolution ODE derived in Shi et al., (2022). Continuous-time analysis of gradient and/or proximal methods themselves form a rich body of the literature. Schropp and Singer, (2000); Alvarez and Attouch, (2001); Alvarez et al., (2002); Attouch et al., (2018) were among the first that formally connects physical systems to gradient algorithms. Principles of classical mechanics Arnold, (1989) have also been applied to this topic Attouch et al., (2016); Wibisono et al., (2016); Wilson et al., (2021). Also, Zhang et al., (2018) shows that refined discretization in the theorem of numeric ODEs scheme achieves acceleration for smooth enough functions. Till now, the finest correspondence between continuous-time ODE and discrete-time Nesterov acceleration is find in Shi et al., (2022). In this paper, we show that NSA easily recovers this high-resolution ODE.
Remark 2.
The correspondence between (24) and (25) underscores another contribution of our work. Proposition 2 fills a gap left open in Shi et al., (2022) – we establish the continuous-time convergence rate for the high-resolution ODE (25) in the convex case, which does not require strong convexity.
@@ 5 NSA for Composite Objectives
4:   $\mathbf{x}_{k+1}^{\prime}=\text{{prox}}_{\eta,h}\left(\mathbf{y}_{k}-\eta\nabla f(\mathbf{y}_{k})\right)$. /* $\eta$ is the step size. */
5:   $\mathbf{x}_{k+1}^{\prime\prime}=\text{{prox}}_{\eta,h}\left(\mathbf{x}_{k}-\eta\nabla f(\mathbf{x}_{k})\right)$.
6:   $\mathbf{x}_{k+1}=\begin{cases}\mathbf{x}_{k+1}^{\prime},&\text{ if }F(\mathbf{x}_{k+1}^{\prime})\leq F(\mathbf{x}_{k+1}^{\prime\prime})\\ \mathbf{x}_{k+1}^{\prime\prime},&\text{ otherwise. }\end{cases}$
7:   $\mathbf{z}_{k+1}=\mathbf{z}_{k}-\frac{\eta}{\alpha_{k}}\nabla f(\mathbf{y}_{k})-\frac{\eta}{\alpha_{k}}\mathbf{g}_{k+1}^{\prime}$, where $\eta\mathbf{g}_{k+1}^{\prime}=\mathbf{y}_{k}-\eta\nabla f(\mathbf{y}_{k})-\mathbf{x}_{k+1}^{\prime}$
@@
By the algorithm update rule, we know $\mathbf{x}_{k+1}^{\prime\prime}$ is a prox-gradient step starting from $\mathbf{x}_{k}$. By the descent property (e.g., Chapter 10 in Beck, (2017)), we know
@@
$F(\mathbf{x}_{k+1}^{\prime})=$ $\;f(\mathbf{x}_{k+1}^{\prime})+h(\mathbf{x}_{k+1}^{\prime})=f(\mathbf{y}_{k}-\eta\nabla f(\mathbf{y}_{k})-\eta\mathbf{g}_{k+1})+h(\mathbf{x}_{k+1}^{\prime})$
@@
$\Delta_{k+1}=$ $\;\frac{\gamma_{k}}{2}\|\mathbf{z}_{k+1}-\mathbf{x}^{*}\|^{2}+f(\mathbf{x}_{k+1})+h(\mathbf{x}_{k+1})-f(\mathbf{x}^{*})-h(\mathbf{x}^{*})$
@@
$+\frac{\eta^{2}\gamma_{k+1}}{2\alpha_{k}^{2}}\|\nabla f(\mathbf{y}_{t})+\mathbf{g}_{k+1}^{\prime}\|^{2}+F(\mathbf{x}_{k+1}^{\prime})-F(\mathbf{x}^{*})$
@@
[…]
+$\mathbf{g}_{k+1}^{\prime},\alpha_{k}\mathbf{x}^{*}+(1-\alpha_{k})\mathbf{x}_{k}-\mathbf{x}_{k+1}^{\prime}-\eta\nabla f(\mathbf{y}_{k})-\eta\mathbf{g}_{k+1}^{\prime}\right>-F(\mathbf{x}^{*}),$
@@
+$\mathbf{g}_{k+1}^{\prime},\alpha_{k}\mathbf{x}^{*}+(1-\alpha_{k})\mathbf{x}_{k}-\mathbf{x}_{k+1}^{\prime}\right>-F(\mathbf{x}^{*})$

### v5

@@ 3 The NSA Algorithm with Inexact Gradient Oracle
Recently, there has been a growing demand for zeroth-order optimization techniques, primarily driven by the need for fine-tuning in contemporary deep learning (e.g., [10.5555/3692070.3694514]). Additionally, many learning scenarios provide access solely to zeroth-order information of the objective function [larson2019derivative]. In such cases, we can estimate the gradient first [flaxman2005online, nesterov2017random, 10.1093/imaiai/iaad014, 10.1093/imaiai/iaac027] and then apply gradient algorithms to solve the problem. Such gradient estimators can be abstractly summarized as an inexact gradient oracle, defined as follows.
@@
Such gradient oracle $\mathcal{G}$ can be guaranteed by zeroth-order information. Formal statements for these properties are in Theorems 2 and 3. Theorem 2 has been proved in the literature by different authors [flaxman2005online, nesterov2017random, 10.1093/imaiai/iaad014]. The key message is that inexact gradient oracles that satisfy conditions (H1) and (H2) can be derived using zeroth-order information. A simple method that ensures conditions (H1) and (H2) is performing finite differences along each coordinate to derive the gradient estimator. Specifically, we consider the most straightforward finite difference gradient estimator defined as follows:
@@
where $\mathbf{e}_{i}$ is the $i$-th unit vector in a fixed canonical basis. The straightforward estimator in (9) serves as an example that fulfills conditions (H1) and (H2). It is important to note that this estimator is entirely deterministic, making $\xi$ a mere placeholder, and thus $\mathbb{E}_{\xi}$ acts as an identity operator. However, (9) is not the only simple estimator that meets these criteria; the estimator presented in [10.1093/imaiai/iaad014] also satisfies conditions (H1) and (H2).
Theorem 2 ([flaxman2005online, nesterov2017random, 10.1093/imaiai/iaad014]).
@@
Consider a function $f\in\mathscr{F}$, where $\mathscr{F}$ denotes Nesterov’s $\mathscr{F}$ functions [nest2003, nest2018]. If at any $\mathbf{x}\in\mathbb{R}^{n}$, we can access the function value $f(\mathbf{x})$, then there exists an oracle $\mathcal{G}$ such that (1) $\mathcal{G}$ satisfies property (H2), and (2) $\mathcal{G}$ can be explicitly calculated using zeroth-order information.
We now present a version of NSA that operates with the inexact gradient oracle, derived from zeroth-order information. The complete procedure is outlined in Algorithm 2.
@@
[…]
@@
$\liminf_{k\to\infty}\beta^{k}\|\nabla f(\mathbf{x}_{k})\|=0\;\text{ for some $\beta>1$},$ (12)
@@
Let $f\in\mathscr{F}_{L}^{1,1}(\mathbb{R}^{n})$ and let $\eta\in(0,\frac{3}{4L}]$. Let $\alpha_{k}=\frac{p}{k+p}$ for some $p>0$. Let $\mathbf{x}_{k}$ be governed by the NSA algorithm (Algorithm 2). If the gradient estimation oracle $\mathcal{G}$ satisfies properties (H1) and (H2), then there exists an error sequence $\{\epsilon_{k}\}_{k\in\mathbb{N}}$, such that $\sum_{k=1}^{\infty}\frac{\epsilon_{k}}{\alpha_{k}}<\infty$, and for all $k=0,1,2,\cdots$, it holds that
@@
$\frac{3}{4}\|\mathcal{G}(f,\mathbf{y}_{k},\epsilon^{\prime};\xi_{k}^{\prime})\|^{2}\leq\nabla f(\mathbf{y}_{k})^{\top}\mathcal{G}(f,\mathbf{y}_{k},\epsilon^{\prime};\xi_{k}^{\prime})\leq\frac{5}{4}\|\mathcal{G}(f,\mathbf{y}_{k},\epsilon^{\prime};\xi_{k}^{\prime})\|^{2},$ (13)
@@
$\frac{3}{4}\|\mathbf{g}_{k}(\mathbf{y}_{k})\|^{2}\leq\nabla f(\mathbf{y}_{k})^{\top}\mathbf{g}_{k}(\mathbf{y}_{k})\leq\frac{5}{4}\|\mathbf{g}_{k}(\mathbf{y}_{k})\|^{2},\quad a.s.$
@@
$\leq$ $\;f(\mathbf{y}_{k})+\left(-\frac{5}{2}\eta+2L\eta^{2}\right)\|\mathbf{g}_{k}(\mathbf{y}_{k})\|^{2},$
where the last inequality follows from (13). Since $\eta\in(0,\frac{3}{4L}]$, we know
@@
$f(\mathbf{x}_{k+1}^{\prime\prime})\leq f(\mathbf{x}_{k})+\nabla f(\mathbf{x}_{k})^{\top}\left(\mathbf{x}_{k+1}^{\prime\prime}-\mathbf{x}_{k}\right)+\frac{L}{2}\|\mathbf{x}_{k+1}^{\prime\prime}-\mathbf{x}_{k}\|^{2}\leq f(\mathbf{x}_{k})-\eta\|\mathbf{g}_{k}(\mathbf{x}_{k})\|^{2}.$ (15)
@@
$\leq$ $\;\sum_{\kappa\in\mathbb{Q}_{+}}\mathbb{P}\left(\liminf_{k\to\infty}k^{2}\log k\log\log k(f(\mathbf{x}_{k})-f^{*})\geq\kappa\right)=0.$
@@ 4 NSA for Nonsmooth Objective
In a Hilbert space $\mathcal{H}$ with inner product $\left$$\cdot$$\cdot\right$> and norm $\|\cdot\|$, with respect to proper lower-semicontinuous convex function $\Psi:\mathcal{H}\to\mathbb{R}$ and step size $\eta$, the Moreau–Yosida proximal operator is defined as a map from $\mathcal{H}$ to $\mathcal{H}$ such that
@@
$\begin{split}\mathbf{x}_{k+1}=&\;\text{{prox}}_{s,f}\left((1-\alpha_{k})\mathbf{x}_{k}+\alpha_{k}\mathbf{z}_{k}\right)\\ \mathbf{z}_{k+1}=&\;\mathbf{z}_{k}+\frac{\eta}{s\alpha_{k}}\left(\mathbf{x}_{k+1}-\big((1-\alpha_{k})\mathbf{x}_{k}+\alpha_{k}\mathbf{z}_{k}\big)\right),\end{split}$ (21)
@@
Let $\partial f(\mathbf{x})$ be the set of subgradients of $f$ at $\mathbf{x}$. By the optimality condition for subgradients (e.g., [RockWets98]), there exists $\mathbf{g}_{k+1}\in\partial f(\mathbf{x}_{k+1})$, such that
@@
$=$ $\;\frac{\gamma_{k+1}}{2}\left\|\mathbf{z}_{k}-\mathbf{x}^{*}\right\|^{2}+\left<\mathbf{g}_{k+1},\alpha_{k}\mathbf{x}^{*}+(1-\alpha_{k})\mathbf{x}_{k}-s\mathbf{g}_{k+1}-\mathbf{x}_{k+1}\right$>
@@
$\leq$ $\;\frac{\gamma_{k+1}}{2}\left\|\mathbf{z}_{k}-\mathbf{x}^{*}\right\|^{2}+\left<\mathbf{g}_{k+1},\alpha_{k}\mathbf{x}^{*}+(1-\alpha_{k})\mathbf{x}_{k}-\mathbf{x}_{k+1}\right>+f(\mathbf{x}_{k+1})-f(\mathbf{x}^{*})$ (since $s\geq\frac{\eta}{2}$)
@@
where the last two lines use convexity of the function $f$. Rearranging terms gives that $\frac{\Delta_{k+1}}{\gamma_{k+1}}\leq\frac{\Delta_{k}}{\gamma_{k}}$. Therefore,
$\frac{\Delta_{k+1}}{\gamma_{k+1}}\leq\frac{\Delta_{k}}{\gamma_{k}}\cdots\leq\frac{\Delta_{1}}{\gamma_{1}},$
@@ 4.1 A Continuous-time Perspective
In this section, we provide a continuous-time analysis for a proximal version of the NSA algorithm (21). In order for us to find a proper continuous-time trajectory of the discrete-time algorithm, Taylor approximation in Hilbert spaces will be used. Therefore, we want the objective function to be smooth. Otherwise, the error terms can get out-of-control and the correspondence between the discrete-time algorithm and continuous-time trajectory may break. After an ODE is derived for the smooth case via Taylor approximation, we could analogously write out an differential inclusion for the nonsmooth case, with derivatives defined in a subdifferential sense (e.g., [doi:10.1137/15M1046095]).
@@
Let $\{\mathbf{x}_{k}\}_{k\in\mathbb{N}}$ and $\{\mathbf{z}_{k}\}_{k\in\mathbb{N}}$ be associated with two curves $X(t)$ and $Z(t)$ ($t\geq 0$). For the first equation in (22), we introduce a correspondence between discrete-time index $k$ and continuous time index $t$: $t=k\sqrt{\eta}$. With this we have $X(t)\approx\mathbf{x}_{k}$, $X(t+\sqrt{\eta})\approx\mathbf{x}_{k+1}$, and $\alpha_{k}=\frac{p}{k+p}=\frac{\sqrt{\eta}p}{t+\sqrt{\eta}p}=\frac{\sqrt{\eta}p}{t}+o(\sqrt{\eta})$. Henceforth, $X$ and $X(t)$ (respectively, $\dot{X}$ and $\dot{X}(t)$) are used interchangeably where the meaning is clear. It holds that
@@
and when $s=o(1)$, we Taylor approximate the trajectories in the Hilbert space $\mathcal{H}$ [cartan2017differential]
@@
$\leq$ $\;p^{2}\left<Z-\mathbf{x}^{*},\dot{Z}\right>+2t\left<\nabla f(X),X-\mathbf{x}^{*}\right>+t^{2}\left<\nabla f(X),\dot{X}\right$> (by convexity)
$=$ $\;p\left<Z-\mathbf{x}^{*},-t\nabla f(X)\right>+2t\left<\nabla f(X),X-\mathbf{x}^{*}\right$>
@@ 4.1.1 A Single Second-order ODE for (22)
Note that the single second-order ODE recovers the high-resolution ODE derived in [2018Understanding]. Continuous-time analysis of gradient and/or proximal methods themselves form a rich body of the literature. [doi:10.1080/01630560008816971, alvarez2001inertial, ALVAREZ2002747, attouch2018fast] were among the first that formally connects physical systems to gradient algorithms. Principles of classical mechanics [arnold1989mathematical] have also been applied to this topic [ATTOUCH20165734, doi:10.1073/pnas.1614734113, JMLR:v22:20-195]. Also, [NEURIPS2018_44968aec] shows that refined discretization in the theorem of numeric ODEs scheme achieves acceleration for smooth enough functions. Till now, the finest correspondence between continuous-time ODE and discrete-time Nesterov acceleration is find in [2018Understanding]. In this paper, we show that NSA easily recovers this high-resolution ODE.
@@ 5 NSA for Composite Objectives
[…]
@@
By the algorithm update rule, we know $\mathbf{x}_{k+1}^{\prime\prime}$ is a prox-gradient step starting from $\mathbf{x}_{k}$. By the descent property (e.g., Chapter 10 in [beck2017first]), we know
@@
$F(\mathbf{x}_{k+1}^{\prime})=$ $\;f(\mathbf{x}_{k+1}^{\prime})+h(\mathbf{x}_{k+1}^{\prime})=f(\mathbf{y}_{k}-\eta\nabla f(\mathbf{y}_{k})-\eta\mathbf{g}_{k+1}^{\prime})+h(\mathbf{x}_{k+1}^{\prime})$
@@
$\Delta_{k+1}=$ $\;\frac{\gamma_{k+1}}{2}\|\mathbf{z}_{k+1}-\mathbf{x}^{*}\|^{2}+f(\mathbf{x}_{k+1})+h(\mathbf{x}_{k+1})-f(\mathbf{x}^{*})-h(\mathbf{x}^{*})$
@@
$+\frac{\eta^{2}\gamma_{k+1}}{2\alpha_{k}^{2}}\|\nabla f(\mathbf{y}_{k})+\mathbf{g}_{k+1}^{\prime}\|^{2}+F(\mathbf{x}_{k+1}^{\prime})-F(\mathbf{x}^{*})$
@@
${\leq}$ $\;\frac{\gamma_{k+1}}{2}\left\|\mathbf{z}_{k}-\mathbf{x}^{*}\right\|^{2}+f(\alpha_{k}\mathbf{x}^{*}+(1-\alpha_{k})\mathbf{x}_{k})+h(\mathbf{x}_{k+1}^{\prime})+\eta\left<\nabla f(\mathbf{y}_{k})+\mathbf{g}_{k+1}^{\prime},\mathbf{g}_{k+1}^{\prime}\right$>
$+\left<\mathbf{g}_{k+1}^{\prime},\alpha_{k}\mathbf{x}^{*}+(1-\alpha_{k})\mathbf{x}_{k}-\mathbf{x}_{k+1}^{\prime}-\eta\nabla f(\mathbf{y}_{k})-\eta\mathbf{g}_{k+1}^{\prime}\right>-F(\mathbf{x}^{*}),$
@@
$+\left<\mathbf{g}_{k+1}^{\prime},\alpha_{k}\mathbf{x}^{*}+(1-\alpha_{k})\mathbf{x}_{k}-\mathbf{x}_{k+1}^{\prime}\right>-F(\mathbf{x}^{*})$